OpenStudy (dejmccray):

evelyn's made a cake in the shape of a circus tent to use as the centerpeice at the celebration. As shown in the illustration, the cake consisted of a cylinder and a cone. The height of the cone is half the height of the cylinder. The volume of the cylinder is 1470 incubed. The diameter of the cylinder is 12 in. What is the volume of the circus tent cake? Use 3014 for pi. Round to the nearest hundredth. Show you work

OpenStudy (dejmccray):

3.14

OpenStudy (dejmccray):

@sammixboo if ur not busy ;/ could u help me?

OpenStudy (anonymous):

Use the volume and diameter of the cylinder to find the height, then use half of that height to find the volume of the cone and add it to 1470. Make sense?

OpenStudy (dejmccray):

okay so 1470= pi*36*h I dont know how to find the height tho

OpenStudy (dejmccray):

@Helloimjohn1234

OpenStudy (anonymous):

You just solve for h like you would any other variable, you need to isolate it on one side

OpenStudy (dejmccray):

okay so what i did was pi*36=113.04 then i divided 1470 by 113.04= 13.004 but thats not right

OpenStudy (anonymous):

If you use 3.14 instead of the pi value on a calculator, then 13.004 is correct. If you do use the pi button that stores all of the digits, then its 12.99765

OpenStudy (dejmccray):

okay but what do i do after 13.004?

OpenStudy (dejmccray):

sorry im not very good at math

OpenStudy (anonymous):

So that is the height of the cylinder, divide it by 2 and use that value to find the volume of the cone using the formula, then add that volume to 1470

OpenStudy (anonymous):

Note that the cone and the cylinder have the same diameter

OpenStudy (dejmccray):

1470=1/3(pi*36)h= 1/3(113.04 or 56.52)h 1/3*113.04= 37.68 1/3*56.52=18.84 which one?

OpenStudy (dulljackel09):

Yea it aint happening XD way to early for this Mental abuse to human= Math

OpenStudy (dejmccray):

lol thanks anyway

OpenStudy (dejmccray):

@iGreen can u help me?

OpenStudy (dejmccray):

@xo.minnie.xox

OpenStudy (anonymous):

Hm..what Dull said xD. Sorry idk

OpenStudy (dejmccray):

aw okay I just need som type of help this is due in a few hours ~.~

OpenStudy (anonymous):

https://www.mathway.com/ Try?

OpenStudy (dejmccray):

@bohotness can u help?

OpenStudy (dejmccray):

@inowalst can u help

OpenStudy (dejmccray):

@johnp12? @Teddyiswatshecallsme?

OpenStudy (dejmccray):

@thcsbabygirl1224

OpenStudy (dejmccray):

@linn99123

OpenStudy (dejmccray):

@Deadpool117

OpenStudy (dejmccray):

@Godlovesme

OpenStudy (anonymous):

i know some one helpfull his name is @Allenismyname

OpenStudy (dejmccray):

okay thanks! @Allenismyname do u think u could help me?

OpenStudy (anonymous):

I'm not sure what problem you are having the information given above seems very helpful.

OpenStudy (dejmccray):

i dont knwo how to find that height and i dont get it @Allenismyname

OpenStudy (godlovesme):

sorry i have no idea :c but i'll tag ppls to help u @Data_LG2 @EclipsedStar @pooja195 @sleepyjess